Session: Oct 2017 (session 3)
Wavelength: 18 cm
To give feedback on this experiment, click
(orfei at Wed Nov 01 16:32:56 2017) Success:
(Jonathan Quick at Wed Nov 01 16:58:55 2017) Success:
No known problems.
(Bo Xia at Wed Nov 01 17:17:19 2017) Success:
(Slepkova, Lysakova, Zadrutsky at Wed Nov 01 18:07:11 2017) Success:
(Kiselev P. at Wed Nov 01 18:08:03 2017) Minor failures: LATE
Scans no0001-no0006 missed
(Slobojaninova N. at Wed Nov 01 18:08:22 2017) Success:
(Uwe at Thu Nov 02 11:20:06 2017) Success:
(Pawel Wolak at Mon Nov 06 09:14:36 2017) Success: RFI
(Jun Yang at Mon Nov 06 10:53:34 2017) Minor failures: WIND
A problem with Onsala Astro-Flexbuff OS disk was found. To solve the problem during the day time, we stopped the observations shortly at 12:30 UT. This problem should have no impact on the recording after checking the output of scan_check. Due to the strong winds, the 25m radio telescope was stowed during 14:05 - 14:55 UT.
(Blaauw at Thu Nov 09 10:28:32 2017) Minor failures:
We got ERROR db -200 dbbc09 Not configured.
We used: form=astro and mk5b_mode=ext,0x00000303,,16.000
The DBBC09 error is caused by a wrong mark5 mask.